# Keyframer: Empowering Animation Design using Large Language Models

### v1

@@ 3.2. Design Goals for AI-Supported Animation Tooling
Overall, our interviews revealed that animation creators were optimistic about using generative AI tools for quickly prototyping animation concepts. Following the challenges and opportunities we identified from the formative interviews, we developed the following design goals (DG) for an AI-supported animation tool:
DG1: Support exploration for animations. Animation creators imagined leveraging the power of AI to generate initial designs and to rapidly explore alternatives. We made these needs part of our system by leveraging LLMs to generate initial starting points users can iterate on. We also took advantage of the flexibility of LLMs to allow users to generate multiple versions of a design at once to aid comparison and selection.
@@
DG3: Empower non-experts to work with animation code. We aimed to reduce the translation burden of turning designs mockups into implemented animations by leveraging the code generation power of LLMs, which we thought could be especially helpful for designers less familiar with code. Additionally, using natural language prompts to generate animation might open up opportunities for non-experts to get started with animation, which led us to expand our audience for testing Keyframer to both professionals and beginners.
@@ 4. Keyframer System
Keyframer is an LLM-powered application we developed for creating animations from static images. Leveraging the code-generation capabilities of LLMs, along with the semantic structure of Static Vector Graphics (SVGs), Keyframer generates CSS for animating SVGs based on user-supplied natural language prompts. Keyframer was designed to support both 1) exploration by enabling users to request, compare, and explore design variants; and 2) refinement, as users are able to iterate on their designs either through reprompting or by editing the LLM-generated animations directly through several designed editor modes. Keyframer is implemented using Open AI’s Chat Completion API  [^10] https://platform.openai.com/docs/guides/text-generation/chat-completions-api, utilizing GPT-4 as its base model. In this section, we describe the design of the Keyframer interface and how its features align with our Design Goals (DGs) described in Section 3.2.
Input. Our system offers an input area where users can paste the code for an SVG image they want to animate. SVG is a standard and popular image format used commonly in illustration due to its scalability and compatibility on many platforms; its use for animations is also established in user interface design and advertising (Drasner, 2017; Head, 2016). In Keyframer, a rendering of the SVG is displayed alongside the code editor so users can preview the visual design of their image as shown in Figure 2. Because SVGs are XML-based images, the LLM is able to use descriptions of the scene embedded in the code, such as object identifiers. For example, in Figure 2, the SVG code for the Saturn illustration contains identifiers such as sky, halos, and clouds. Object identifiers are commonly defined by designers by naming layers when creating graphics in applications like Illustrator or Sketch.
@@
GPT Prompt. Following DG3, our system allows users to create animations by entering a natural language prompt into a text input field. As an open-ended text field, users can make a request for a single design (“Make the planet spin”) or for multiple design variants (“Create 3 designs where the stars twinkle”, DG1). The user clicks the Generate Animations button to begin their request.
Before passing a user request to GPT, we supplement their prompt with the full raw SVG XML and specify the format of the LLM response. For example, we request that the response consists solely of CSS, as generating both SVG and CSS led to longer response times and higher risk of exceeding token limitations. Further, we ask for each CSS snippet to be accompanied by a descriptive explanation that can help users inspect what the LLM generated design is intended to look like. The full details of our prompt can be found in Appendix A.5.
GPT Output. Once a prompt request has started, we stream the response from GPT, which consists of one or more CSS snippets as shown in Figure 3. The streaming response provides feedback to the user about whether the request is still in progress, while also allowing users to preview the LLM’s response. Following DG3, we support users unfamiliar with CSS to interpret animation code by prompting the LLM to generates descriptive names for keyframes. Generated names provide a hint for the design that will be rendered, such as ‘flame-flicker’ or ‘flash-sparkles’. Each time we detect a complete snippet, we render the animation inline as described in the next section.
Figure 3. GPT Output section where the streamed response from the LLM is displayed in advance of rendering each CSS snippet.
GPT Output section where streamed response from GPT is displayed in advance of rendering each CSS snippet.
Figure 3. GPT Output section where the streamed response from the LLM is displayed in advance of rendering each CSS snippet.
Rendering. The Rendering section consists of, 1) a visual rendering of each animation along with 1-sentence explanation generated by the LLM, and 2) a series of editors for modifying the design.
@@
Once a code snippet is detected (e.g., the delimiter between code snippets is found in the GPT-output), we inject the CSS snippet along with the original SVG code onto the page to display the design. As we render multiple designs with the same underlying SVG on a single page, we have the LLM provide a unique class for each CSS snippet (e.g., .design-1, .design-2, etc.). Visual renderings appear side by side to invite comparison and aid users in selecting the design that best matches their vision.
@@
Figure 4. Rendering section for viewing generated designs side by side and editing response code in the Code or Properties editors.
@@
Rendering section for viewing generated designs side by side and editing response code in Code or Properties editor.
Figure 4. Rendering section for viewing generated designs side by side and editing response code in the Code or Properties editors.
@@
To provide granular controls for editing the generated animations (DG2), we offer two kinds of editors in Keyframer. When a user clicks on a given design, the editors below display the CSS snippet in two formats: a Code Editor where the CSS can be edited directly, and a Properties Editor, a dynamically created UI-layer for editing CSS properties. Following DG3, the Properties Editor is designed to be approachable to users less familiar or comfortable with editing CSS syntax and is modeled after UI elements one might see in graphics editors like Illustrator.
@@
The Code Editor is implemented with CodeMirror and has syntax highlighting and autocompletion for CSS. We re-inject the revised CSS onto the page on every keystroke so that users can preview their changes immediately. The Properties Editor provides property-specific UI for editing the code (DG3); for example, for editing colors, we provide a color picker and for editing timing function curves, we provide a dropdown of defaults (like linear and ease-in), along with a bezier editor for entering custom timing curves. Animation properties and keyframes are grouped together by identifier to improve readability. Figure 5 shows how an example code snippet appears in both the code editor and the properties editor. Edits in either mode are immediately reflected in the companion mode to support direct manipulation (Kery et al., 2020).
@@
Figure 5. The Code and Properties Editors enable users to edit generated animation code directly. The editors are bi-directional so that any edits in one editor are reflected in the corresponding editor. The Properties editor adds a UI layer for editing animation properties without having to understand CSS syntax.
@@
Figure 5. The Code and Properties Editors enable users to edit generated animation code directly. The editors are bi-directional so that any edits in one editor are reflected in the corresponding editor. The Properties editor adds a UI layer for editing animation properties without having to understand CSS syntax.
Iteration. To support user exploration in the animation creation process (DG1), we offer a feature to allow users to iteratively build on a generated animation using prompts. Underneath every generated design is a button + Add New Prompt; clicking this button opens up a new form at the bottom of the page for a user to extend their design with a new prompt. In this way, users can continually add additional prompts in sequence to iterate on their animation. When a user adds a new iteration, our prompt requests the LLM to extend existing CSS code as outlined in A.5.
Saved Designs Sidebar & Summary. We enable users to star designs to favorite them and add them to a sidebar, as displayed on the right-hand side of Figure 6. Clicking on a given design in the sidebar will scroll to where the design was generated so the user can revisit its corresponding and code. We also have a Summary mode which hides all text editors and displays the animations along with their prompts, enabling users to quickly revisit their previous prompts and designs.
@@ 5. User Study Methodology
We conducted a user study to answer the following research questions about Keyframer:
@@ 5.1. Study Procedures
To understand the user experience of designing animations with Keyframer, we facilitated 90-minute sessions over video conference in which individuals built a set of animations with the tool.
At the beginning of the study, we presented a hypothetical scenario in which the participant is asked to help a friend, who is a children’s book illustrator, design animations for her personal website using Keyframer. We chose this framing as it assumes the illustrations have already been created, enabling us to focus the user study solely on animation. We briefly introduced Keyframer to the participants, demonstrating how to generate animations from SVG images using natural language prompts, how to generate multiple designs at once, and how to use the Code and the Properties Editors. We also introduced participants to a third-party system called Boxy [^11] https://boxy-svg.com that could be used during the study to explore the provided SVG images and read their element identifiers to use in prompting (see A.3 for further details).
@@
We then started the main part of the study by asking participants to create animations using two provided images, displayed in Figure 7. Activity 1 involved generating animations for a header image of the website, and Activity 2 consisted of animating an image thanking newsletter subscribers. Both SVG images were prepared by the authors of this paper in advance (full details on the creation of these SVGs can be found in A.2). For each activity, participants were first asked to explore the SVG using Boxy and brainstorm how they would like to animate the elements in the image. Then the participants used Keyframer freely for 15 minutes to generate animations with the goal of creating designs they would be happy to share with their friend. We asked participants to think aloud as they crafted their prompts and responded to LLM output. At the end of both activities, we interviewed participants about their experience creating animations using Keyframer. All participants completed Activity 1 followed by Activity 2.
After both activities, we asked participants a series of Likert scale questions on their satisfaction with the animations they created in 15 minutes, the helpfulness of the system for generating ideas for animations, and the helpfulness of the system for refining animations. Participants were also asked open-ended questions about their prompting strategies, feature requests, and potential use cases of using Keyframer for future work. The full details of our study design, research instruments, and study format can be found in A.4.
@@ 5.2. Participant Recruitment
To recruit participants, we advertised our user study on an location-based, general-purpose Slack channel with 2,500 members internal to a large technology company. We invited anyone interested in creating animations and described the study as centering around a ‘Generative AI design tool for prototyping animations,’ specifying that no coding experience was required. In addition, we reached out via email to participants in our formative study for snowball sampling among their co-workers.
Those interested in participating first filled out a screener survey, answering several questions about their job title, gender identity (for recruitment balancing purposes), any prior experience they might have with animation or coding, and whether they had experience using any AI tools. From the total pool of respondents, we selected a subset representing a range of programming and animation experience, as several professional animators interviewed in our formative study pointed out that Keyframer might be especially helpful for novices.
We mapped out a 2x2 matrix representing animation and programming experience such that we had representation from four groups: High Code and High Animation (HCHA), Low Code and Low Animation (LCLA), High Code Low Animation (HCLA), and Low Code High Animation (LCHA). A participant was considered low code if they had little or no prior programming experience; in contrast, a high code participant writes software professionally. Low animation users had little or no experience with animation design, while those categorized as high animation included users who had taken classes in animation design in college, do animation work professionally, or have taught animation coursework. We selected participants to have a balanced number of participants from each group.
@@ 5.3. Data Collection & Analysis
We recorded the video conference sessions (audio and video), giving us 19.5 hours of video for analysis. Additionally, we instrumented the application with logging features to create timestamped activity logs, tracking the prompts users entered, the full response from the LLM to each request, and which editors they used to edit the LLM output, if any. The anonymized logs were stored locally in the participant’s browser and exported and shared with the researchers by the participants at the conclusion of each session. Transcripts were auto-generated by the video conferencing software and reviewed and edited by the research team.
Following the procedures of thematic analysis (Braun and Clarke, 2006), two researchers met weekly over the course of two months to analyze the transcripts and logs from our sessions. We inductively coded the transcripts for participants’ experience using Keyframer for animation design. In addition to analyzing the interview transcripts, we quantitatively analyzed all the prompts users generated in both activities, all the animations and code generated, and whether or not users edited the output from the LLM (and if applicable, what editor mode they used for their edits), comparing across all four groups.
Additionally, we again followed the procedures of thematic analysis to analyze prompting styles and semantic prompt types. We first independently coded a subset of 25 prompts (or 12% of all unique prompts) over multiple rounds to create a set of prompting styles and semantic prompt types. After reviewing and discussing any discrepancies in our application of these codes, we then split the list of prompts and independently coded half, discussing until we reached agreement on all prompts.
@@ 6. User Study Results
From 54 total survey respondents, we selected 13 participants (6 female, 7 male) to try out Keyframer in our user studies. Table 1 summarizes our participants, with three participants in each skill level category (with the exception of the HCHA group that had four participants, including one who was also in our formative study). Participants had a range of job titles including brand content manager, research engineer, and UX/UI designer, with a mix of skill levels in both animation and programming. Four participants (EP1, EP2, EP3, EP17) had no prior programming experience. All reported having tried out existing AI tools (11 out of 13 have used ChatGPT, and 7 have used Dall·E and/or Midjourney), largely to test their capabilities and for fun.
@@
Figure 8. Six frames taken from animations generated by EP4 (HCLA) and EP9 (HCHA). The Activity 1 Saturn animation from EP4 has the sparkles fade in and out independently, the clouds fade in, the halos fade in one after another, and the specks alternate colors from yellow to orange to pink. The Activity 2 Rocketship animation from EP9 has the rocketship move up and down, the clouds grow and shrink in size, and items in the background (the moon and specks) shift down to give the appearance of the rocketship lifting off.
@@
Overall, participants were satisfied by their experience with Keyframer. When asked to rate their satisfaction with the animations they generated in only 15 minutes, participants on average rated $3.93.9$ (SD = $1.01.0$), which is between ‘satisfied’ (4) and ‘neutral’ (3). Participants generated 223 designs (94 for Activity 1, 129 for Activity 2). On average, each participant generated 17.2 designs (SD = $5.55.5$). An example of two participants’ final animations are displayed in Figure 8.
Participants were happily surprised by the efficacy of the system to turn natural language instruction into animations. Keyframer supported participants with limited expertise in motion design to generate animations from scratch efficiently. For instance, EP6 stated, “This is just so magical because I have no hope of doing such animations manually…I would find [it] very difficult to even know where to start with getting it to do this without this tool.”
Participants appreciated that the system sped up their animation prototyping process: “I think this was much faster than a lot of things I’ve done… I think doing something like this before would have just taken hours to do” (EP2). Those without experience in programming animations found Keyframer “democratize[d] this whole creative process” and “highlights creativity more so than the mechanics of [the code],” letting users focus on higher level goals instead of having to know exactly how an animation could be implemented in CSS. Seeing LLM-generated CSS in Keyframer also offered novices “an interesting little window to see what’s happening behind the scenes” (EP1) to connect code to animations.
Even professional motion designer EP13 saw the potential of Keyframer to extend his capabilities: “Part of me is kind of worried about these tools replacing jobs, because the potential is so high. But I think learning about them and using them as an animator — it’s just another tool in our toolbox. It’s only going to improve our skills. It’s really exciting stuff.”
System Performance. On average, $90.490.4$ (SD = $36.736.7$) lines of code were generated by the LLM in the final designs for Activity 1, and $87.287.2$ (SD = $63.563.5$) for Activity 2. The generated CSS from GPT-4 included CSS properties such as opacity, fill color, visibility, scale, and timing-function (a full list of generated CSS properties is listed in A.6). Keyframer was able to generate code for animations in a timely manner, taking on average $17.417.4$ seconds (SD = $9.79.7$s, max = $62.062.0$s, min = $5.95.9$s) to generate a code instance from a prompt. In addition, the code generated by the LLM were generally high-quality. Only 6.76.76.76.7% of GPT-4 generated CSS (n=15) resulted in code with syntactical errors (the details of which are provided in A.7). In those cases, participants would regenerate code using the same prompt or edit the prompt.
The most common challenges users encountered when requesting animations were having the LLM correctly interpret group versus individual level behavior, and getting the LLM to correctly execute time-sensitive sequencing. For example, in Activity 1, a common prompt request was to ‘animate the sparkles to twinkle.’ Users almost always meant for each of the three sparkles to animate independently, but because there was a group-level specification in the SVG (<g id="sparkles">), the LLM applied the animation to the three sparkles as a group. Users would then reprompt to specify independent animation, e.g., ‘animate each sparkle to twinkle.’ Another common challenge was getting two elements to move together, especially in the rocketship illustration. While this could be accomplished by grouping items in the underlying SVG code, doing so solely in CSS is more difficult as it requires matching time delays and duration, which the LLM did not always successfully execute.

### v2

@@ 3.2. Design Goals for AI-Supported Animation Tooling
Overall, our interviews revealed that animation creators were optimistic about using generative AI tools for quickly prototyping animation concepts, while still wanting ultimate control over the final design. Based on our findings from our formative interviews, we developed the following design goals (DG) for an AI-supported animation tool:
DG1: Support exploration for animations. Animation creators imagined leveraging the power of AI to generate initial designs and to rapidly explore alternatives. Our system supports this goal by leveraging LLMs to generate initial starting points users can iterate on. We also take advantage of the flexibility of LLMs to allow users to generate multiple versions of a design at once to aid comparison and selection.
@@
DG3: Empower non-experts to work with animation code. We aimed to reduce the translation burden of turning designs mockups into implemented animations by leveraging the code generation power of LLMs, which we thought could be especially helpful for designers less familiar with code. Additionally, using natural language prompts to generate animations might open up opportunities for non-experts to get started with animation, which led us to expand our audience for testing Keyframer to both professionals and beginners.
@@ 4. Keyframer System
Keyframer is an LLM-powered application for creating animations from static images. Leveraging the code-generation capabilities of LLMs, along with the semantic structure of Static Vector Graphics (SVGs), Keyframer generates CSS for animating images based on user-supplied natural language prompts. Keyframer supports 1) exploration by enabling users to request and compare design variants and 2) refinement, as users can iterate on their designs either through reprompting or by editing the LLM-generated animations directly. Keyframer is implemented using GPT-4.
4.1. Animating an image with code generation
With Keyframer, user provide an SVG image to animate, enter natural language prompts to generate CSS code, and can preview the resulting animations rendered from the generated code.
The Input area is where users can paste the code for an SVG image they want to animate as shown in Figure 2. SVG is a standard and popular image format used commonly in illustration and user interface animations(Drasner, 2017; Head, 2016). Because SVGs are XML-based images, the LLM is able to use descriptions of the scene embedded in the code, such as object identifiers. For example, in Figure 2, the SVG code for the Saturn illustration contains identifiers such as sky, halos, and clouds. Object identifiers are commonly defined by designers by naming layers when creating graphics in applications like Illustrator or Sketch.
@@ 4.1. Animating an image with code generation
Following DG3, our system allows users to create animations by entering a natural language user prompt into a text input field. Users can make a request for a single design (“Make the planet spin”) or for multiple design variants (“Create 3 designs where the stars twinkle”, DG1). Before passing a user request to GPT, we supplement their prompt with the full SVG XML and specify the output format, instructing the LLM to return CSS animation code and ask for each code snippet to be accompanied by a descriptive explanation [^14] The full details of our prompt can be found in Appendix A.5. We chose to use CSS rather than Javascript as CSS enables a wide range of animation effects and can be more interpretable for beginners (DG3).
Once a request has been submitted, we stream the GPT Output, which consists of one or more CSS snippets as shown under “Keyframer Output” in Figure 1. Following DG3, our prompt asks the LLM to generates descriptive names for keyframes to make the code more interpretable for beginners (for example, adding CSS keyframe names like‘flame-flicker’ or ‘flash-sparkles’).
@@
Each time we detect a complete snippet, we render a preview of the animation inline. Once a code snippet is detected, we inject the CSS snippet along with the original SVG code onto the page to display the animation. As we render multiple designs with the same underlying SVG on a single page, we have the LLM provide a unique class for each CSS snippet (e.g., .design-1, .design-2, etc.). When generating multiple design variants, visual renderings appear side by side to invite comparison and aid users in selecting the design that best matches their vision. For example, in Figure 3 where the user prompts “generate 3 designs where the sky fades into different colors,” the LLM generates three distinct options to spark ideas.
@@
Figure 3. Rendering section for viewing generated designs side by side and editing output code in the Code or Properties editors.
[…]
@@
Rendering section for viewing generated designs side by side and editing output code in Code or Properties editor.
4.2. Editing generated animations
@@
Keyframer provides two editors for users to directly modify generated animation code (DG2): a Code Editor where the CSS for a given design can be edited directly, and a Properties Editor, which presents dynamically created UI for editing CSS properties. Following DG3, the Properties Editor is designed to be approachable to users less familiar with code and is modeled after UI elements one might see in graphics editors like Illustrator.
@@
The Code Editor is implemented with CodeMirror and has syntax highlighting and autocompletion for CSS. We re-inject the revised CSS onto the page on every keystroke so that users can preview their changes immediately. The Properties Editor provides property-specific UI for editing the code (DG3) such as a color picker and a dropdown and bezier editor for timing functions. Figure 4 shows how an example code snippet appears in both the code editor and the properties editor. Edits in either mode are immediately reflected in the companion mode to support direct manipulation (Kery et al., 2020).
@@
Figure 4. The Code and Properties Editors enable users to edit generated animation code directly. The editors are bi-directional so that any edits in one editor are reflected in the corresponding editor. The Properties editor adds a UI layer for editing animation properties without having to edit code.
[…]
@@
4.3. Iterating on animations
To support user exploration (DG1), Keyframer allows users to iteratively build on a generated animation using additional prompts. Underneath every generated design is a + Add New Prompt button; which lets users extend their design with a new prompt. In this way, users can continually add additional prompts in sequence to iterate on their animation. When a user adds a new iteration, our prompt requests the LLM to modify or extend existing CSS code as outlined in A.5. To support review, users can bookmark their favorite animations or turn on a Summary mode to hide all text editors, allowing them to quickly revisit previous prompts and designs as shown in Figure 5.
@@ 5. User Study Methodology
Our user study was designed to answer the following research questions:
@@ 5.1. Study Procedures
We facilitated 90-minute sessions over video conference in which individuals built a set of animations with Keyframer. Participants were presented with a hypothetical scenario in which they are asked to help a children’s book illustrator design animations for her personal website, using her illustrations. We began by briefly demoing Keyframer, demonstrating how to enter prompts, request multiple designs, and use the editors to edit generated animations. We also introduced a third-party system called Boxy [^15] https://boxy-svg.com that could be used to explore the provided SVG images and read their element identifiers to use in prompting (see A.3 for further details).
@@
We then started the main part of the study by asking participants to open Keyframer in their browser and create animations using two provided images (Saturn and Rocketship), displayed in Figure 6. Both SVG images were prepared by the authors of this paper in advance (full details on the creation of these SVGs can be found in A.2). For each activity, participants were first asked to explore the SVG using Boxy and brainstorm ideas for animating the image. Then the participants used Keyframer freely for 15 minutes with the goal of creating designs they would be happy to share with the illustrator. We asked participants to think aloud as they crafted their prompts and responded to LLM output. All participants completed Activity 1 followed by Activity 2.
After both activities, we used a semi-structured protocol to ask participants open-ended questions about their prompting strategies, feature requests, and potential use cases of using Keyframer for future work. The full details of our study design, research instruments, and study format can be found in A.4.
@@ 5.2. Participant Recruitment
We advertised our user study on a message channel with 2,500 members internal to a large technology company. In addition, we reached out via email to participants in our formative study for snowball sampling among their co-workers.
Those interested in participating first filled out a screener survey, answering several questions about their job title, gender identity (for recruitment balancing purposes), any prior experience with animation or coding, and whether they had experience using AI tools. From the total pool of respondents, we selected a subset representing a range of programming and animation experience, as several professional animators interviewed in our formative study pointed out that Keyframer might be especially helpful for novices. We mapped out a 2x2 matrix representing animation and programming experience such that we had representation from four groups: High Code and High Animation (HCHA), Low Code and Low Animation (LCLA), High Code Low Animation (HCLA), and Low Code High Animation (LCHA). A participant was considered low code if they had little or no prior programming experience; in contrast, a high code participant writes software professionally. Low animation users had little or no experience with animation design, while those categorized as high animation included users who had taken classes in animation design in college or do animation work professionally.
@@ 5.3. Data Collection & Analysis
We collected screen recordings and audio from each video conference session, giving us 19.5 hours of video for analysis. Additionally, we instrumented the application with an anonymized, session-based logging features to create timestamped activity logs tracking the prompts users entered, the LLM response for each request, and which editors participants used to edit the LLM output, if any. Transcripts auto-generated by the video conferencing software were reviewed and edited by the research team and after each session, researchers wrote reflective memos to capture salient moments. Our data collection and study protocol were reviewed and approved by a research ethics committee and legal counsel at a large technology company.
Following the procedures of thematic analysis (Braun and Clarke, 2006), two researchers met weekly over the course of two months to analyze the transcripts and logs from our sessions to inductively identify themes around the Keyframer user experience. We also quantitatively analyzed the prompts users wrote (which were coded blind of participant identifiers), the animations and code generated, and user editing patterns in modifying generated animation code, comparing across all four groups.
We identified a set of prompting styles by first having two authors independently code a subset of 25 user prompts (or 12% of all unique prompts) over multiple rounds. After reviewing and discussing any discrepancies in our application of these codes, the same two researchers split the list of prompts and independently coded half, discussing until we reached agreement on all prompts.
@@ 6. User Study Results
Study Participants From 54 survey respondents, we selected 13 participants (6 female, 7 male), with 3 or 4 participants in each animation and code skill level category (Table 1) to gather qualitative feedback from a diverse range of perspectives. Those with ‘high animation’ experience have worked on advertising campaigns, data visualization, and UI animations using After Effects (EP8, EP9, EP10, EP12, EP13) and Javascript and CSS (EP7, EP9, EP10). Four participants (EP1, EP2, EP3, EP13) had no prior programming experience. All reported having tried out existing AI tools (11 out of 13 have used ChatGPT, and 7 have used Dall·E and/or Midjourney), largely to test their capabilities and for fun.
@@
Figure 7. Five frames taken from animations generated by EP4 (HCLA) and EP9 (HCHA). The Activity 1 Saturn animation from EP4 has the sparkles fade in and out independently, the clouds fade in, the halos fade in one after another, and the specks alternate colors from yellow to orange to pink. The Activity 2 Rocketship animation from EP9 has the rocketship move up and down, the clouds grow and shrink in size, and items in the background (the moon and specks) shift down to give the appearance of the rocketship lifting off.
@@
General Impressions of Keyframer Participants were happily surprised by the efficacy of the system to turn natural language instruction into animations. Beginners without animation experience found it easy to get started with Keyframer: “This is just so magical because I have no hope of doing such animations manually…I would find [it] very difficult to even know where to start with getting it to do this without this tool” (EP6). Similarly, EP5 shared how Keyframer “democratize[d] this whole creative process” and “highlights creativity more so than the mechanics of [the code],” letting users focus on higher level goals instead of having to know CSS syntax. Participants appreciated that the system sped up their animation prototyping process: “Doing something like this before would have just taken hours” (EP2).
Even professional motion designer EP13 saw the potential of Keyframer to complement his animation process, sharing how LLMs are “just another tool in our toolbox. It’s only going to improve our skills.”
System Performance. On average, $90.4$ (SD = $36.7$) lines of code were generated by the LLM in the final designs for Activity 1, and $87.2$ (SD = $63.5$) for Activity 2. The generated CSS from GPT-4 included CSS properties such as opacity, fill color, visibility, scale, and timing-function (a full list of generated CSS properties is listed in A.6). Keyframer generated animation code in a timely manner, taking on average $17.4$ seconds (SD = $9.7$s, max = $62.0$s, min = $5.9$s) to generate a code instance. In addition, the code generated by the LLM were generally high-quality, with only 6.76.7% (n=15) of code snippets with syntactical errors (the details of which are provided in A.7). An example of two participants’ final animations are displayed in Figure 7.